# Towards Next-Generation Urban Decision Support Systems through AI-Powered Generation of Scientific Ontology using Large Language Models - A Case in Optimizing Intermodal Freight Transportation

### v1

@@ 3.2.2 LLM-powered Ontology Generation
    Build Concept Taxonomy: A step focuses on organizing and structuring the identified concepts into a hierarchical framework. Building a concept taxonomy involves arranging concepts (or classes) identified during the conceptualization phase into a taxonomic (hierarchical) structure based on ”is-a” relationships. This hierarchy represents a superclass-subclass relationship between broader and more specific concepts.
@@
    Define Ad Hoc Relationship: This step is a crucial phase in the ontology development process. It involves identifying and specifying the relationships that exist between the concepts that are specific to the domain being modeled and do not necessarily fit into well-defined hierarchical (is-a) or associative (part-of) relationships.
After creation, the scenario-based ontologies generated by our proposed methods undergo validation against the documentation of existing domain ontologies, which typically describe high-level, generic domain knowledge. This validation is conducted using the reasoning capabilities of LLMs. The purpose of this process is to ensure accurate and relevant identification of domain concepts, as well as logical mapping of relationships between concepts and classes. Additionally, this procedure ensures semantic consistency across the AI-created ontology through a semantic mapping process that leverages the existing domain ontology. After these procedures, the finalized ontology output is then stored in an OWL format, which is interoperable and compatible with popular ontology management software tools and modern semantic web technologies.
3.2.3 The ontology implementation
@@
A specific use case involves converting the LLM-generated OWL ontology into SPARQL (SPARQL Protocol and RDF Query Language) using Protégé. SPARQL is primarily used to query and manipulate data stored in the Resource Description Framework (RDF) format (Kumar and Kumar, 2013). This capability enables decision support applications to more effectively discover, retrieve, and integrate data resources from RDF databases. In this setting, the ontology can be used to facilitate data management from relational database (Lubbad, 2018), enabling an data provision pipeline to support data-driven insights during the development of urban decision support system.
3.3 Qualitative Result Evaluation
(a) Outer Level
(b) Inner Level
Figure 4: We show a human-created ontology, which defines a subclass-of relationship. This relationship, expressed as “A is-a B,” signifies that A represents a more specific or specialized category than B. In essence, this arrangement delineates the hierarchical structuring of knowledge.
We evaluate the feasibility of our methodology through a comparative analysis, juxtaposing our AI-generated ontology with the widely recognized Pizza Ontology. Class materials [^2] Available at: http://www.meteck.org/swdsont.html and the work discussed by Rector et al. (2004), depict the challenges faced in the domain of web ontology language description and generation logic. These challenges often can lead to conceptual misunderstandings and underscore the necessity for expert domain knowledge to effectively structure, comprehend, and ultimately construct knowledge representations.
In the realm of the use of an automated ontology creation process, we qualitatively verified its proper generation, which contrasts with the Pizza representation, to ensure the LLM-generated ontology have valid . We focus on two sections of this ontology; under Figure 4, we present the human-created ontology, which shows an “is-a” or “subclass-of” relationship. This subtype or subclass, indicates a specialization-generalization relationship.
@@ 3.3 Qualitative Result Evaluation
It’s crucial to highlight the current empirical evidence that underscores the robust capabilities of applying methodologies akin to those outlined in Fernández-López et al. (1997), now seamlessly executed through Large Language Models (LLMs). The ontology illustrated in Figure 4 amalgamates inputs from undisclosed sources, revealing the logical interconnections among the presented classes and subclasses. Conversely, the data utilized for the automated pizza ontology stems from a distinct singular source, presented in a plain format.
The modular components outlined in the subsections of the ontology pizza source document delineate a systematic breakdown of information, subsequently processed through the proposed framework and LLM. Figures 5(b) and 5(a) delineate pertinent sections and mirror the same logical components and properties showcased in the diagram.
(a) Outer Level
(b) Inner Level
Figure 5: An unprocessed automated ontology per section fully generated via the LLM. This ontology depicts vegetable toppings embedded as individuals in the LLM generated model. The inclusion of individual instances composed by artichokes, mushrooms, onion tomatoes resembles the human made ontology.
Under 5(b) we represented a narrow display of detail level represented via a pizza class and corresponding sub classes that carry styles, characteristic tool and components. In addition, label under ’Pizza’ a deeper level can be seen. In 5(a) This more general level of understanding about the context of pizza
4 Case Study: Ontology in Intermodal Freight Transportation
@@
In this section, we present a compelling use case illustrating the application of our methodology in tackling a real-world decision optimization challenge. Our focus lies in the intricate realm of intermodal freight transportation across the United States, where we harness the power of integrating multiple datasets and simulations. The primary objective of the study is to optimize this complex system with the goals of enhancing fuel efficiency and, at the same time, reducing Greenhouse Gas (GHG) emissions and total operation costs. This endeavor stands as a demonstration to the capability of our methodology in addressing multifaceted, real-world challenges with tangible environmental and economic implications in urban systems.
@@ 4.1 Data Sources and Simulation Models
Our methodology integrates three primary data sources alongside external data from academic sources, enriching the knowledge and analysis of intermodal transportation networks as represented under Figure 3. The Freight Analysis Framework (FAF) (U.S. Department of Transportation, 2024) and its accompanying manual serve as foundational resources, identifying crucial freight transportation hubs across diverse regions. These hubs provide a framework for mapping significant geographic and logistic parameters that define major nodes within transportation networks. Similarly, the FTOT (U.S. Department of Energy, 2024b) and its manual are instrumental in permitting the routing engine to perform calculations under the criteria of the shortest path; this transportation network is represented via shapefiles. The generated ontology will permit LLMs to perform calculations and handle services such as an endpoint REST API to compute unique routes for various nodes, encompassing road, rail, and water modes, while also delineating the properties and relationships within these networks. Manuals ensure a comprehensive understanding of the methodologies and algorithms applied in these tools, enhancing the accuracy of simulations and optimizations in our model.
The Alternative Fueling Stations (AFS-BTS) dataset (U.S. Department of Energy, 2024a) averages 1,521 locations per state and encompasses a total of 80,601 locations across the US providing valuable insights into alternative fueling stations. Of particular interest are the 69,314 electric charging stations, highlighting electric vehicles as the most prevalent renewable fuel type. The dataset also includes information on fuel types including Electric, E85, Liquefied Petroleum Gas, Bio-Diesel, Compressed Natural Gas, Renewable Diesel, Liquefied Natural Gas, and Hydrogen. This dataset facilitates detailed modeling of CO2 emissions, a critical aspect of our optimization and emission reduction strategies. We incorporate a fourth data source. Academic contributions, such as methodological approaches for calculating equivalent CO2 emissions, among other sources, enhance our understanding and integration of emissions considerations within the LLM framework, further refining our strategies for optimization and sustainability in intermodal freight transportation.
In addition to the aforementioned primary data and simulations encompassing operational costs, fuel consumption, and GHG emissions across various modes of freight transportation in the US, our approach integrates a comprehensive array of open GIS datasets. These datasets delineate critical fuel and transportation infrastructure locations, leveraging resources such as the Homeland Infrastructure Foundation-Level Data (HIFLD) for fuel station and supply locations, and the OpenStreetMap network for vital highways and rail systems. Furthermore, to capture the dynamic physical processes inherent in intermodal freight transportation, we incorporate real-time traffic congestion and accident information. This invaluable data is sourced from real-time traffic sensor feeds accessible via web APIs. Noteworthy examples include the Radar Detection System operated by the Tennessee Department of Transportation, along with the WAZE API.
4.2 Ontology-driven Decision Optimization
The methodology, empowered by LLM, initially utilized for creating the pizza ontology, is now repurposed to generate a scenario-based knowledge representation. This representation integrates various intermodal freight transportation data and simulation outputs, alongside other urban datasets. The resulting real-time decision support workflow targets the optimization of urban logistic systems across extensive geographic areas, demonstrating the versatility and efficacy of the LLM-powered approach, as depicted in Figure 6.

### v3

@@ 4.2.3 LLM-powered Ontology Construction - Knowledge Extraction
    Build Concept Taxonomy: A step focuses on organizing and structuring the identified concepts into a hierarchical framework. Building a concept taxonomy involves arranging concepts (or classes) identified during the conceptualization phase into a taxonomic (hierarchical) structure based on “is-a” relationships. This hierarchy represents a superclass-subclass relationship between broader and more specific concepts.
@@
    Define Ad Hoc Relationship: This step is a crucial phase in the ontology development process. It involves identifying and specifying the relationships that exist between the concepts that are specific to the domain being modeled and do not necessarily fit into well-defined hierarchical (is-a), associative (part-of) relationships, or attribute (has property) relationships.
After creation, the scenario-based ontologies generated by our proposed methods undergo validation against the documentation of existing domain ontologies, which typically describe high-level, generic domain knowledge. This validation is conducted using the reasoning capabilities of LLMs. This process aims to ensure accurate and relevant identification of domain concepts, as well as logical mapping of relationships between concepts and classes. Additionally, this procedure ensures semantic consistency across the AI-created ontology through a semantic mapping process that leverages the existing domain ontology. After these procedures, the finalized ontology output is then stored in an OWL format, which is interoperable and compatible with popular ontology management software tools and modern semantic web technologies.
@@
A specific use case involves converting the LLM-generated OWL ontology into SPARQL using Protégé. SPARQL is primarily used to query and manipulate data stored in the RDF format (Kumar and Kumar, 2013). This capability enables decision support applications to more effectively discover, retrieve, and integrate data resources from RDF databases. In this setting, the ontology can be used to facilitate data management from relational database (Lubbad, 2018), enabling a data provision pipeline to support data-driven insights during the development of urban decision support systems.
4.2.4 Ontology Evaluation
@@
It is crucial to highlight the current empirical evidence that underscores the robust capabilities of applying methodologies akin to those outlined in Fernández-López et al. (1997), now seamlessly executed through Large Language Models (LLMs). The human-crafted ontology amalgamates inputs from undisclosed sources, revealing the logical interconnections among the presented classes and subclasses. Conversely, the data utilized for the automated pizza ontology stems from a distinct singular source, presented in a text format through the PDF files. The modular components in pizza source document provide a systematic breakdown of information, which is then processed through the proposed framework and LLM. The automated pizza ontology, showed decomposed in an outer 111 Available at https://github.com/jtupayachi/Auto_Ontology/blob/main/figures/SECTION1_PIZZA.png and inner  [^2] Available at https://github.com/jtupayachi/Auto_Ontology/blob/main/figures/ONTOLOGY_TOPPINGS.png sections mirror the logical components and properties resulting from the method.
Ontologies, define a subclass-of relationship. This relationship, could be expressed as “A is-a B,” signifies A represents a more specific or specialized category than B. In essence, this arrangement delineates the hierarchical structuring of knowledge. We evaluate the feasibility of our methodology through a comparative analysis, juxtaposing our AI-generated ontology with the widely recognized Pizza Ontology (Human-crafted). The work discussed by Rector et al. (2004) depicts the challenges faced in the domain of web ontology language description and generation logic. These challenges often can lead to conceptual misunderstandings and underscore the necessity for expert domain knowledge to effectively structure, comprehend, and ultimately construct knowledge representations. The use of an automated ontology creation process requires qualitatively verifying its proper generation, which contrasts with the Pizza representation, to ensure the LLM-generated ontology is valid. The human-created ontology  [^3] Available at: https://github.com/jtupayachi/Auto_Ontology/blob/main/comparison_pizza/pizza.owl, shows an “is-a” and “subclass-of” relationships. This subtype or subclass indicates a specialization-generalization relationship. An automated ontology  [^4] Available at: https://github.com/jtupayachi/Auto_Ontology/blob/main/comparison_pizza/automated.rdf per section is fully generated via the LLM. This ontology depicts vegetable toppings embedded as individuals in the LLM-generated model. The inclusion of individual instances composed of artichokes, mushrooms, onions, and tomatoes resembles the human-made ontology. The automated ontology modules generated (one per each text section - topic) were merged using the OWL API.
Similarly, we represented the inner components of human-crafted ontology  [^5] Available at: https://github.com/jtupayachi/Auto_Ontology/blob/main/figures/onto_outer.png, through the main pizza class and its corresponding subclasses, which encompass styles, characteristics, and components. Additionally, a deeper level of the labeling under ’Pizza’ 666 Available at: https://github.com/jtupayachi/Auto_Ontology/blob/main/figures/onto_inner.png provides a broader context and a more general understanding of this topic. Several methods have been proposed to evaluate ontology outcomes and assess the alignment of frameworks and methods with the intended topic. Redmond and Noy (2011) argue that direct textual comparison becomes ineffective in one-to-one comparisons. They focus instead on verifying and mapping axioms and target ontologies using a differential engine built on the OWL API, which calculates structural differences between ontologies. Similar approaches are followed by other researchers, such as Malone et al. (2010) with the Bubastis tool and OWLDiff by Kremen et al. (2011). It is important to note that these methods are particularly suited for scenarios where ontologies undergo modifications, aiming to uncover such differences effectively.
A more suitable approach for comparing our framework with a human-crafted ontology involves using CQs Kommineni et al. (2024). The approach utilizes pre-defined CQs, such as ”Write the competency questions based on the abstract level concepts for describing …”, to evaluate the utility of an ontology. Qualitative comparisons between ontologies created using different methods can also be conducted by examining the ontology’s capability to address individual CQs. Noy and Musen (2002) proposed a CQs-based method that focuses on merging and aligning ontologies, where the authors highlight alignment points such as ”Industrial org” and ”Commercial Organization”. Although these terms differ, they represent similar concepts. This method relies on user inputs, which, when combined with the PROMPT tool (Noy and Musen, 2003), can be used to compare and evaluate the structure within each ontology. The evaluation considers the differences in the Prompt-Diff table, specifically examining the number of rows added, deleted, or changed in their operation or mapping-level columns. Other studies have also introduced various approaches for evaluating the creation of ontologies using CQs, similarly as used in Kommineni et al. (2024).
Table 1: Qualitative pizza ontology comparison using CQs and qualitative analysis. A SPARQL platform is employed for the execution of the queries and interface with the (.rdf or .owl) files. The keyword “ontology” in this table references the url of the uploaded ontological files.
CQ Queries on Human-Crafted Ontology Queries on Automated Ontology
What are the primary categories within the pizza ontology? PREFIX: rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> SELECT DISTINCT ?class WHERE { ?class rdf:type owl:Class . } Pizza, PizzaBase, Food, Spiciness 105 in total PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> SELECT DISTINCT ?class WHERE { ?class rdf:type ontology:Class .} Food, Process, Business, Culture 34 in total
What are the main ingredients of a pizza? Not found in ontology Not found in ontology PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>, PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#> SELECT ?individual ?label WHERE { ?individual rdf:type ontology:Ingredients . OPTIONAL { ?individual rdfs:label ?label }} Cheese, Mozzarella, Tomatoes 18 in total
What are the different variations of toppings used for a pizza? PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>, PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#> SELECT ?class WHERE { ?class rdfs:subClassOf ontology:VegetableTopping . } TomatoTopping, HotGreenPepperTopping, JalapenoPepperTopping, ArtichokeTopping, AsparagusTopping, OnionTopping, PeperonataTopping, CaperTopping, OliveTopping 22 in total PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>, PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#> SELECT ?subclass WHERE { ?subclass rdfs:subClassOf ontology:PizzaToppings .} CheeseToppings, FruitToppings, HerbAndSpiceToppings, MeatToppings, OtherToppings, SauceToppings, SeafoodToppings, VegetableToppings
Which dishes use the dough as a main ingredient? Not found in ontology Not found in ontology PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> SELECT DISTINCT ?dish WHERE { ?dish rdf:type ontology:Dish . ?dough rdf:type ontology:Dough .} Calzone, Pizza
What are the Vegetable Toppings? PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>, SELECT ?individual WHERE { ?individual rdf:type ontology:VegetableToppings . } Tomato, Hot Green Pepper, Jalapeno Pepper, Artichoke, Asparagus, Onion 22 in total PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> SELECT ?vegetableTopping WHERE { ?vegetableTopping rdf:type ontology:VegetableToppings .} Artichokes, Mushrooms, Onion, Tomatoes
What are the different types of Pizzas? PREFIX rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#>, PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#> SELECT ?pizzaType ?label WHERE { ?pizzaType rdf:type owl:Class . ?pizzaType rdfs:subClassOf* ontology:Pizza . OPTIONAL { ?pizzaType rdfs:label ?label . } AmericanHot, Cajun, Capricciosa, Caprina, Fiorentina, FourSeasons 25 in total Not found in ontology Not found in ontology
4.2.5 CQ-based Qualitative Analysis
Unlike many existing studies that focus on generating foundational domain ontologies and knowledge representations (Kommineni et al., 2024), our method expands existing domain ontologies to create scenario-based ontologies from new datasets or simulations, facilitating practical decision support solutions. To evaluate the AI-generated ontology from our method, we adopt the commonly accepted ontology evaluation methods using CQs and qualitative analysis based on justifications from human-expert (Bezerra et al., 2013; Ren et al., 2014). Table 1 presents a side-by-side comparison of the AI- and human-generated pizza ontologies, and it includes the formulation of CQs and the SPARQL queries used for evaluation. The results indicate that the outputs of the AI-generated ontology are sensible as with those from the human-crafted ontology.
In our scenario, it is important to note that the primary data source for creating the pizza ontology is not identical to the one used for the human-crafted ontology. This discrepancy may lead to variations in scope, classes, and subclasses, as well as the resulting SPARQL queries. To perform a refined evaluation, it is crucial for the initial sources of both approaches to be identical. As shown in Table 1, the content from the AI-generated ontology is queried using SPARQL and provides logical and reasonable responses to address each CQ. Although there may be some differences in the actual content, the outputs demonstrate that the classification and alignment of the AI-generated pizza ontology are sensible, logical, and practical. Our comparative analysis ensures that the AI-generated ontology can accurately reflect knowledge without providing hallucinated or fabricated results.
5 Case Study: Optimizing the Intermodal Freight Transportation System
@@
In this section, we present a compelling case study to demonstrate the application of our methodology in addressing a real-world decision optimization challenge. Our focus is on the analysis of intermodal freight transportation across the United States, where we leverage the integration of multiple datasets. The primary objective of the study is to optimize this complex system with the goals of enhancing fuel efficiency and, at the same time, reducing Greenhouse Gas (GHG) emissions and total operation costs. This endeavor stands as a demonstration of the capability of our methodology in addressing multifaceted, real-world challenges with tangible environmental and economic implications in urban systems.
Figure 3: The demonstration includes prompt templates used in the LLM-powered workflow for constructing ontology and knowledge, as well as the design of the data provisioning system.
@@ 5.1 Data Sources and Simulation Models
Our methodology integrates three primary data sources (FAF and FTOT) alongside an academic source this enriches the knowledge and analysis of intermodal transportation networks as represented under Figure 2. The Freight Analysis Framework (FAF) (U.S. Department of Transportation, 2024a) and its accompanying manual serve as foundational resources, identifying crucial freight transportation hubs across diverse regions. Identified hubs provide a framework for mapping significant geographic and logistic parameters that define major nodes within transportation networks. Similarly, the FTOT (U.S. Department of Transportation, 2024b) and documentation are instrumental in permitting the creation of the intermodal network. Calculations under the criteria of the shortest path; this transportation network is represented via shapefiles. The generated ontology will allow a decision support program, script, or LLM, to perform calculations and handle services such as an endpoint REST API to compute unique routes for various nodes, encompassing road, rail, and water modes, while delineating the properties and relationships within these networks. Manuals ensure a comprehensive understanding of the methodologies and algorithms applied in these tools, enhancing the accuracy of simulations and optimizations in our model. Academic contributions, such as methodological approaches for calculating equivalent CO2 emissions, among other sources, enhance our understanding and integration of emissions considerations within the LLM framework, further refining our strategies for optimization and sustainability in intermodal freight transportation. In addition to the aforementioned primary data and simulations encompassing operational costs, fuel consumption, and GHG emissions across various modes of freight transportation in the US, our approach integrates the comprehensive array of intermodal GIS transportation networks provided by FTOT. The location of major critical transportation hubs is documented through FAF.
5.2 Ontology and Knowledge Graph Construction
We applied our LLM-based methods to the three documents of the Freight Analysis Framework (FAF) and the Freight and Transportation Optimization Tool (FTOT), deriving an OWL ontology. This ontology was then converted into a knowledge graph in RDF format to support more practical data integration and management applications.
Figure 4: Demonstration of prompt templates used to instruct the LLM.
Figure 3 illustrates the implementation details, including knowledge representations and database components. The instructions and prompt templates provided to the LLM for generating the ontology at each critical step are demonstrated in Figure 4. From the supplied PDF files, we perform the cleaning phase, which includes stop word removal, punctuation, and special character handling. Next, the process relies on glossary generation, a crucial step where the text is simplified into a set of relevant words. This step is supported by an iterative validation process that enhances robustness and filters out words that do not persist through multiple iterations. We supply the elaborated prompt with the title and keywords of the provided section and request the LLM for confirmation on the content it provides. After the completion of this iterative process, the LLM moves to the next function, which verifies consistency and ensures that the relevant words glossary are aligned with the section. This glossary serves as the input for taxonomy generation, where we establish the relationships and hierarchical order within the glossary of words and concepts. Our OWL file generator then transforms the identified classes, subclasses, and individuals into a readable OWL file. Following this, an iterative process ensures the LLM finalizes the output until the stop condition in the metadata is met. While challenges may arise, different methods can be applied to address them. The LLM is then used to transform the OWL output and construct a query for the creation of a table schema, later represented as the entity-relationship diagram (ERD). More technical details on how the ERD is implemented using the AI-generated ontology are provided in Section 5.4. The AI-generated intermodal ontologies is composed of three separate ontology modules:
  1. 1.
    The first module is the ontology supporting the FAF data, providing essential data descriptions for developing a data provisioning system. This system encompasses factors such as types of transportation, cargo deliveries, geographical areas, and data sources for transportation data.
  2. 2.
    The second module includes the FTOT network, which incorporates the spatial data (GIS data) necessary for calculating routes and inter-hub distances.
  3. 3.
    The third module comprises the problem statement and optimization models derived from various academic resources. This module focuses on greenhouse gas emission calculations, optimization formulations, and intermodal transportation agents.
To formulate the decision support workflow and problem-solving strategies, we use the third ontology module to derive two attributes of the core concepts involved in optimizing intermodal freight transportation. These attributes include ’properties,’ which indicate the connections and relationships between different concepts, and ’individuals,’ which are specific instances created for particular objects or entities to define core decision support metrics. Demonstration of the AI-generated ontology and sample attributes are provided in Figure 5.
Figure 5: AI-generated ontology for describing concepts, data, and simulation outputs for optimizing intermodal freight transportation.
5.3 Data Provisioning System: Technologies and Implementation
A robust data provisioning system plays a critical role in the discovery, querying, and retrieval of data necessary for supporting informed decisions. In our case study, we utilize an AI-generated ontology and knowledge graph to guide the design and development of our data provisioning system. This system consists of essential databases for storing intermodal transportation data and simulation outputs, as well as web API endpoints for querying and accessing these datasets. Given the nature of the FAF and FTOT data, which encompass large transportation networks (topologically connected graph data) connecting numerous origins and destinations (e.g., intermodal transits and warehouses in major US cities) using different transportation modes and routes, we employ both graph database and a relational database to store the integrated intermodal transportation attributes.
For the graph database management system, we used neo4j to store the FAF transportation network as graph data elements (nodes, edges connecting them, and attributes of nodes and edges). To import the FAF network into a neo4j database, we derived the transportation network based on the Origin-Destination (OD) pairs documented in the FAF data. The resulting network (represented by a graph) describes the simplest form of the connectivity of roads, rails, and rivers across the US, with graph components (e.g., nodes as cities and edges as routes) serving as building blocks to join attributes and metrics necessary for decision optimization. The primary purpose of the neo4j database is to identify all possible combinations of modes and routes (represented by lists of nodes and edges) based on any user-defined OD pairs requested through web APIs. To optimize neo4j’s performance in traversing the network, we store only the transportation network and its topology within the neo4j database, including only essential attributes (e.g., node and edge IDs, modes, and edge distances and slopes). These essential attributes are added to the graph to allow the neo4j database to apply constraints during the network traversal. Attributes related to the route (edge) geometry are pre-calculated using spatial analysis tools and included in the graph as numerical attributes. Additional graph element attributes pertaining to nodes and edges, which characterize different facets of intermodal transportation as decision metrics, are not imported into the graph database. Instead, these non-essential attributes are stored in the relational database. This practice of partitioning non-essential attributes from the graph aims to enhance neo4j’s performance in navigating the large-scale transportation network.
Once all alternative combinations of modes and routes are identified from the neo4j database, we use the lists of node and route IDs for each combination to retrieve the necessary data and simulation outputs as decision metrics from the relational database. In this study, we use PostgreSQL, with its schema guided by the knowledge graph, to integrate and store vast and diverse urban data and simulation outputs (e.g., FTOT and URIs to traffic data APIs) that characterize intermodal transportation processes. Examples of these metrics include total GHG emissions, total shipment time, total fuel consumption, and total operation cost. These metrics are either pre-calculated and stored in the relational database or dynamically calculated using real-time traffic information (e.g., congestion indexes and road closure information) retrieved from third-party data APIs through time-scheduled scripts deployed in a cloud-based server environment. The AI-generated ontology guides the design of the relational database schema and the information searches in the relational databases based on user-defined criteria and requirements (e.g., inclusion or exclusion of GHG emissions) for optimization.
5.4 Data Provisioning System: Database Schema Design
The ontology and its knowledge graph are primarily used to construct the database schema and Entity Relationship Diagram (ERD) for the PostgreSQL database, as well as to provide essential metadata and descriptions for each metric used in decision-making. By traversing the graph stored in neo4j, we can identify all possible routes between a set of user-defined origins and destinations. Based on these identified routes, our ontology facilitates the discovery and retrieval of all required metrics from the PostgreSQL database by mapping individual metrics to their associated graph components in the identified routes, such as edges and nodes. In terms of implementation, our proposed LLM-powered methods can export the ontology of intermodal transportation data and simulations, such as FAF and FTOT, in OWL format. Many existing ontology management software tools provide straightforward methods and plugins for converting OWL to relational database schemas (Mogotlane, 2014). In our case study, we use Protégé and its DataMaster plugin to convert the AI-generated ontology into the PostgreSQL database schema. DataMaster allows users to import database schemas into an ontology and vice versa (Mogotlane and Fonou-Dombeu, 2016). If needed, custom scripts in Python or Java can also be created to translate the ontology and knowledge graph into a relational schema. Libraries such as RDFLib (for Python) or Apache Jena (for Java) can assist with parsing and processing OWL files.
Figure 6: Automated Knowledge Graph Representation
A subset of the ERD for the relational database is depicted in the purple boxes in Figure 3. The sample subset of the FAF network data stored in the neo4j database is shown in Figure 3. Both databases provide web API endpoints as URIs to facilitate information discovery, query, and retrieval through common internet communication protocols, such as REST. The client-side application of a decision support system can send requests to these URIs, which are later manually added to the knowledge graph to enhance data observability. Under Figure 6, we present a snippet of the knowledge graph that includes both the generated pizza knowledge graph and the FAF dataset knowledge graph. This represents an intermediate step between the conceptual ontology and the creation of relational databases. We utilized the OntoGraf plugin from Protégé for visualization.
Table 2: Ontology-guided selection of decision support metrics and workflow using FTOT and FAF ontology.
CQ Automated Ontology Prefix, Query and Output Query Output Description Dataset
What are the primary scenario parameters within the FTOT Ontology? PREFIX: rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> PREFIX: rdfs: <http://www.w3.org/2000/01/rdf-schema#> PREFIX: owl: <http://www.w3.org/2002/07/owl#> SELECT ?parameter WHERE { ?parameter rdf:type base:ScenarioParameters . } Scenario_Name, Scenario_Description, RMP_Commodity_Data, Destinations_Commodity_Data, Disruption_Data, 20 in total The parameters, metadata related parameters, details on raw material producers, destinations, and information on disruption events. FTOT
What are the primary scenario inputs within the FTOT Ontology? PREFIX: rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> PREFIX: rdfs: <http://www.w3.org/2000/01/rdf-schema#> PREFIX: owl: <http://www.w3.org/2002/07/owl#> SELECT ?input WHERE { ?input rdf:type base:ScenarioInputs . } geospatialinformation, networkattributes, facilities, origins, destinations, 10 in total Inputs listed include GIS information, attributes related the network such as traffic or disruptive information, location for facilities, origins and destinations. FTOT
Describe the geographical components of the FAF dataset? PREFIX: rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> PREFIX: rdfs: <http://www.w3.org/2000/01/rdf-schema#> PREFIX: owl: <http://www.w3.org/2002/07/owl#> SELECT ?class ?individual WHERE { ?class rdfs:subClassOf <<---URL--->/ontology#Geography> . ?individual rdf:type ?class . } DomesticOrigin, ForeignOrigin, DomesticDestination, ForeignDestination, 9 in total Geographical components, such as origin and destination. FAF
List the main transportation hubs? PREFIX: rdf: <http://www.w3.org/1999/02/22-rdf-syntax-ns#> PREFIX: rdfs: <http://www.w3.org/2000/01/rdf-schema#> PREFIX: owl: <http://www.w3.org/2002/07/owl#> SELECT ?region WHERE ?region rdf:type ontology:Regions . Mobile-Daphne-Fairhope, Orlando-Deltona-Daytona-Beach, Chicago-Naperville, Tucson-Nogales, St-Louis-St-Charles-Farmington, Los-Angeles-Long-Beach, Philadelphia-Reading-Camden, 24 in total List of individual hubs identified by the SPARQL query on the regions class within the FAF ontology. FAF
5.5 Decision Support Strategies - Formulation
Our ontology-driven decision support workflow aims to optimize urban logistics systems across large geographic areas, showcasing the versatility and effectiveness of the LLM-powered approach. By defining CQs aligned with the optimization goals, SPARQL queries are crafted to navigate the AI-generated knowledge representation. This process aids in problem formulation and the selection of decision support metrics, as illustrated in Table 2. These queries return specific concepts, variable names, and relevant metadata, providing urban researchers with a clearer understanding of the available data and simulation resources. The same approach can also be used to identify constraints from the ontology, offering an intuitive and interactive method for identifying pertinent variables from integrated data systems, thereby supporting the development of effective decision support workflows. The case study presented in this section utilizes a subset of FAF and FTOT data to establish a use case for identifying the optimal combination of shipping modes and routes from Nashville to New Orleans. The conceptual workflow is depicted in Figure 7. Traditional manual approaches to decision optimization often require extensive domain expertise in the dynamics of intermodal freight transportation events, along with data- and simulation-driven insights for quantifying various decision support criteria, such as operational costs and emissions.